The Ferrell Transportation Company uses a responsibility reporting system to measure the performance of its three investment centers: Planes, Taxis, and Limos. Segment performance is measured using a system of responsibility reports and return on investment calculations. The allocation of resources within the company and the segment managers’ bonuses are based in part on the results shown in these reports. Recently, the company was the victim of a computer virus that deleted portions of the company’s accounting records. This was discovered when the current period’s responsibility reports were being prepared. The printout of the actual operating results appeared as follows. Determine the missing pieces of information below. (Round intermediate calculations and final answers to 0 decimal places, e.g. 1,255.) Planes Taxis Limos Service revenue $ $495,000 $ Variable costs 5,490,000 295,000 Contribution margin 245,000 423,500 Controllable fixed costs 1,495,000 Controllable margin 75,000 188,500 Average operating assets 24,950,000 1,450,000 Return on investment 10 % 7 % %
Answer #1
Planes Taxis Limos
sales a)9480000 495000 g)718500
variable cost 5490000 d)250000 295000
Contribution margin b)3990000 245000 423500
controllable fixed cost 1495000 e)170000 h)235000
controllable margin

c)2495000

75000 188500
Average operating asset 24950000 f)1071429 1450000
Return on investment 10% 7% i)13%

C)Controllable margin = Average operating asset*ROI

                 = 24950000 *10%

                 = 2495000

b)Contribution margin =controllable fixed cost+ controllable margin

                    = 1495000 +2495000

                 = 3990000

a)Sales =contribution margin + variable cost

              = 3990000+5490000

              = 9480000

d)Variable cost =sales -contribution margin

           = 495000-245000

           =250000

e)controllable fixed cost =contribution margin -controllable margin

                       = 245000 -75000

                       = 170000

f)Average operating asset =controllable margin /ROI

                             = 75000/.07

                                = 1071429

g)sales = 295000+423500

            = 718500

h)controllable fixed cost = contribution margin - controllable margin

                            = 423500 -188500

                              =235000

i)Return on investment = controllable margin /average operating asset

                        = 188500/1450000

                          = .13 or 13%